Chair: David Smith

Next meeting chair: Ahmad

Attended: Earl, David Schunck, David Smith, Jasper, Mars, Jamie, Eloise, Jay, Lim, Ahmad.

Apologies: Simin

General Issues:

Handover issues:

R41066

Yarragadee 12m

Disk VSN: HOB+0068/8000/1024 | Data volume at beginning: 74550 GB

  • 2100UT Experiment started late (EARL)

AUM051 clkoffset not working - maser had a slow drift during the day, which apparently caused errors, but just a monitoring problem.

Make sure RfS switch is on for all mixed mode sessions at Hb - otherwise in VGOSS mode.

Hobart 12m

Disk VSN: Flexbuff-2hb

  • 1900UT Experiment started okay (Jay)
  • 2252UT At changeover, saw alert about difference between formatter and maser delays (~0.8μs) restarted DBBC (Mars)
  • 2326UT DBBC restarts but issue is still there, was advised to leave it unless it went over ~1.0μs (Mars)
  • 0700UT delay difference has slowly grown throughout shift, now up to ~0.93μs (Mars)
  • 0937UT - dbbc3 restarted because of bad sampler values. Multiple restarts of the dbbc did not work. Checked the pduhb to find that the RfSwOnForS wasn't on. Switched it on and did a power cycle of the dbbc3. Experiment restarted at 1128UT. Data obtained from 1211UT (Prad).
  • NOTE: I don't think RfSwOnForS wasn't on. I checked and made sure it's on at start of exp. and scan recording( 4 stream) confirms that. (Jay)

Katherine 12m

Disk VSN: Flexbuffke

  • 1900UT Experiment started okay (Jay)

Yarragadee 12m

Disk VSN: HOB+1007| Data volume at beginning: 0GB

  • 1900UT Experiment started okay (Jay).

AUM052

Bad sampler at Hb - resync and power cycling worked for about 5 minutes. Might require input swap. Jamie looking into it. Earlier is better for power cycling.

Hobart 12m

Disk VSN: Flexbuff-2hb

  • 1930UT Experiment started okay (Jasper)
  • 2358UT Bad sampler error. reconfigured the dbbc3hb. Affected scans: 246-2358 to 247-0141 (Simin)
  • 1018UT Bad sampler F, dbbc3hb power cycled. Affected scans 246-1018 - 246-1141 (Eloise)
  • 19:00 Sampler went bad, not resolved before experiments end. Affected scans: 247 - 1853 to 247 - 1928 (EARL)

Katherine 12m

Disk VSN: Flexbuffke

  • 1930UT Experiment started okay (Jasper)

Yarragadee 12m

Disk VSN: HOB+1007| Data volume at beginning: ~6900GB

  • 1930UT Experiment started okay (Jasper)

RV155

Yarragadee 12m

Disk VSN: HOB+0101/16000 | Data volume at beginning: 0GB

  • 1745UT Experiment started late. First scan is from 1745UT. (Jay).
  • Observer forgot the shift i think. Got a msg on my mobile about no fs activity on yg. after checking, found out my mobile number was still there in log monitor, which is from last night. highly probable that no one actually observed since my shift (Warren is away to ke, so not sure about who did the day shift as well). Anyway, couldn't resist to quickly set up yg for next experiment (didn't feel right to just leave things). So seems like It was basically me observing all night (Jay).

Everyone make sure you know when your shifts are, and contact the person before you and the person after you.

R41067 Note: Unable to view the experiment module allocation page since VPN access got nuked. Will just have to rely on people at Yg switching modules correctly - though I've had to call on several occasions when they've forgotten, so we should probably get that sorted (Eloise)

Modules page working now.

z22255 mk5=evlbi? about half the data was dropped and not written to disk. Not immediately fixable - 100% CPU utilisation. Don't do data analysis on the flexbuffs (flexbuffhb)

Hb: Last 15 minutes suffered very high data loss (~50%) due to conflicting process (JMc)

It seems that multiple people were also using the flexbuff at the time.

t2155 Minor issues with mk5.

Yarragadee 12m

Disk VSN: HOB+0064/8000/1024 | Data volume at beginning: … GB

  • 1758 UT Experiment started late (missed first 5 scans) due to some issues with mk5 - but disk swap, on-site reboot (spoke w Martin there) and a few re-syncs w ddbc got it sorted. (Earl).

Yg - correlation reports in R1 sessions have noisy results. Jamie is looking into it. Usually due to a phase instability. Jamie investigating maser and a cavity tune to improve q factor of maser. Next R1 will be the first after tuning. Cavity tune might improve things. If anything usual happens, let Jamie know. Keep an eye on Yg.

slogit not working well. Give Jamie a call if there is an issue. Copying the proc file gives setup changes. Slogit should be working now.